OpenStudy (anonymous):

The molar mass of calcium is 40.08, oxygen is 16.0, and hydrogen is 1.01. What is the mass of one mole of calcium hydroxide (Ca(OH)2)?

OpenStudy (anonymous):

Use the formula\[moles = \frac{ mass }{ molar mass }\] and rearrange to get \[mass = moles \times molar mass\] Next you would find the molar mass of Ca(OH)2 which is 40.08 + (16x2) + (1.01x2) = 74.1 Then since we have one mole we just put the numbers into the formula: mass = 1 x 74.1 mass = 74.1g

OpenStudy (jfraser):

add up all the pieces....